# Nonlinear scalar field equations with a critical Hardy potential

### v1

@@ 1. Introduction
Our focus on the critical case $\mu=\mu_{c}$ should be contrasted with the subcritical case $\mu<\mu_{c}$. For subcritical potentials, the quadratic form remains coercive (or can be made so with an equivalent norm), allowing for a more standard application of variational methods within $H^{1}(\mathbb{R}^{N})$ or related weighted spaces. The subcritical case was of interest to many authors (see e.g. [12, 17, 18, 31] and references therein). For instance, Li, Li, and Tang [17] studied the subcritical problem under Berestycki-Lions-type conditions, establishing the existence of a ground state solution within the $H^{1}(\mathbb{R}^{N})$ space. We find that for the critical potential, any nonnegative solution is genuinely singular and does not belong to $H^{1}(\mathbb{R}^{N})$. This highlights a fundamental difference in the qualitative nature of solutions as the potential transitions from subcritical to critical. Other related works, such as those by Deng, Jin, and Peng [6], Felli, Pistoia [8], Smets [23], and Terracini [26], have investigated problems combining the Hardy potential with the critical Sobolev exponent $2^{*}=\frac{2N}{N-2}$, which introduces a different set of compactness challenges related to the nonlinearity itself.
@@
    There exists a nontrivial solution $u_{0}\in X^{1}(\mathbb{R}^{N})$ of (1.2) that satisfies the Pohožaev constraint $\mathcal{M}$}.
@@
Next, we investigate the existence of non-radial solutions. We follow the ideas of [19]. Suppose that $N\geq 4$ and write $\mathbb{R}^{N}=\mathbb{R}^{M}\times\mathbb{R}^{M}\times\mathbb{R}^{N-2M}$, where $2\leq M\leq\frac{N}{2}$. For $x\in\mathbb{R}$ we will write $x=(x_{1},x_{2},x_{3})$, where $x_{1},x_{2}\in\mathbb{R}^{M}$ and $x_{3}\in\mathbb{R}^{N-2M}$. We introduce an action $\tau:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ by
@@
Suppose that (G1)–(G5) are satisfied and that $g$ is odd. Then
$\inf_{{\mathcal{M}}\cap X_{\tau}\cap X^{1}_{\mathcal{O}}(\mathbb{R}^{N})}J\geq 2\inf_{{\mathcal{M}}}J$
@@ 2. Functional setting
Let $u\in X^{1}(\mathbb{R}^{N})\setminus H^{1}(\mathbb{R}^{N})$. Note that, thanks to e.g. the characterization of the norm (2.4) from [27, 28], we know that $\nabla u\in L^{2}(\mathbb{R}^{N}\setminus B(0,\varepsilon))$ for every $\varepsilon>0$. Moreover, since $u\not\in H^{1}(\mathbb{R}^{N})$, $\nabla u\not\in L^{2}(B(0,\varepsilon))$ for every $\varepsilon>0$. Fix $y\in\mathbb{R}^{N}\setminus\{0\}$ and choose $\varepsilon<\frac{|y|}{2}$. Then $\nabla u(\cdot-y)\not\in L^{2}(B(y,\varepsilon))$, but $B(y,\varepsilon)\subset\mathbb{R}^{N}\setminus B(0,\varepsilon)$, so $\nabla u(\cdot-y)\not\in L^{2}(\mathbb{R}^{N}\setminus B(0,\varepsilon))$, meaning that $\nabla u(\cdot-y)\not\in X^{1}(\mathbb{R}^{N})$. ∎
@@ 3. Variational setting
Then, by the deformation lemma [30, Lemma 2.3], we may find a flow $\eta\in{\mathcal{C}}([0,1]\times X^{1}(\mathbb{R}^{N});X^{1}(\mathbb{R}^{N}))$ such that $\eta(1,\gamma(\cdot))\in\Gamma$ and $J(\eta(1,\gamma(t^{*})))\leq c-\varepsilon<c$. Moreover, for every $t\neq t^{*}$, $J(\eta(1,\gamma(t)))\leq J(\gamma(t))<c$. Hence, for every $t\in[0,1]$ we have $J(\eta(1,\gamma(t)))<c$. Thus $\sup_{t\in[0,1]}J(\eta(1,\gamma(t)))<c$ and we obtain a contradiction. ∎
@@
For every $\varepsilon\in\left(0,\frac{c}{2}\right)$ we may choose a path $\gamma\in\Gamma$ such that $\sup_{t\in[0,1]}(J\circ\Phi)(0,\gamma(t))\leq c+\varepsilon$. Hence, by Ekeland’s variational principle [30, Theorem 2.8], there exists $(t,u)\in\mathbb{R}\times X^{1}(\mathbb{R}^{N})$ such that
@@ 4. Profile decomposition of Pohožaev-Palais-Smale sequences
Since $w^{1}\in H^{1}(\mathbb{R}^{N})$ is a weak solution to (4.5), from [22, Theorem 2], there is $\alpha>0$ such that $|w^{1}(x)|\lesssim e^{-\alpha|x|}$. For $x\in B(0,R)$ we have $|x-y_{n}|\geq|y_{n}|-|x|\geq|y_{n}|-R\to\infty$, so one has
@@ 6.2. Regularity of the solution
and $|u|^{p-2}\in L^{N/2}(\Omega)$. Therefore, by Brezis-Kato lemma (see [24, Lemma B.3]), $u\in L^{q}_{\mathrm{loc}}(\Omega)$ for any $q<\infty$. Then, by Caldéron-Zygmund inequality (see [24, Lemma B.2]), $u\in W^{2,q}_{\mathrm{loc}}(\Omega)$ and next, by Sobolev embedding theorem, $u\in C^{1,\alpha}_{\mathrm{loc}}(\Omega)$. Since $x_{0}$ was an arbitrary point, we conclude $u\in C^{1,\alpha}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$. ∎
@@ 6.3. There is a radial solution, which is a minimizer on $\mathcal{M}$}
$\int_{\mathbb{R}^{N}}\frac{|u^{*}|^{2}}{|x|^{2}}\,dx\geq\int_{\mathbb{R}^{N}}\frac{u^{2}}{|x|^{2}},\quad\int_{\mathbb{R}^{N}}G(u^{*})\,dx=\int_{\mathbb{R}^{N}}G(u)\,dx.$
@@ 6.4. Nonnegative solutions do not belong to $H^{1}(\mathbb{R}^{N})$
From [23, Proposition 2.1], $)u\not\in H^{1}(B(0,\widetilde{\rho}$)). ∎
@@ 7. Existence of non-radial solutions
In this section we will prove Theorem 1.3. Following [19], we recall the subspace $X_{\tau}$ given by (1.4) and the group $\mathcal{O}=\mathcal{O}(M)\times\mathcal{O}(M)\times\mathcal{O}(N-2M)$ acting on $\mathbb{R}^{N}=\mathbb{R}^{M}\times\mathbb{R}^{M}\times\mathbb{R}^{N-2M}$. Repeating the reasoning from [19, Corollary 3.2] we see that the following concentration-compactness principle holds in the subspace $H^{s}_{\mathcal{O}}(\mathbb{R}^{N})$ of $\mathcal{O}$}-invariant functions in $H^{s}(\mathbb{R}^{N})$.
@@
Let $R>1$. Now, as in [19, Remark 4.2], we choose an odd and smooth function $\varphi:\mathbb{R}\rightarrow[-1,1]$ such that $\varphi(x)=1$ for $x\geq 1$, and $\varphi(x)=-1$ for $x\leq-1$. Then, as shown in [19, Remark 4.2], the function (with a small abuse of notation $w_{R}(x)=w_{R}(|x|)$)

### v2

@@ 1. Introduction
Our focus on the critical case $\mu=\mu_{c}$ should be contrasted with the subcritical case $\mu<\mu_{c}$. For subcritical potentials, the quadratic form remains coercive (or can be made so with an equivalent norm), allowing for a more standard application of variational methods within $H^{1}(\mathbb{R}^{N})$ or related weighted spaces. The subcritical case was of interest to many authors (see e.g. [12, 17, 30] and references therein). For instance, Li, Li, and Tang [17] studied the subcritical problem under Berestycki-Lions-type conditions, establishing the existence of a ground state solution within the $H^{1}(\mathbb{R}^{N})$ space. We find that for the critical potential, any nonnegative solution is genuinely singular and does not belong to $H^{1}(\mathbb{R}^{N})$. This highlights a fundamental difference in the qualitative nature of solutions as the potential transitions from subcritical to critical. Other related works, such as those by Deng, Jin, and Peng [6], Felli, Pistoia [8], Smets [22], and Terracini [25], have investigated problems combining the Hardy potential with the critical Sobolev exponent $2^{*}=\frac{2N}{N-2}$, which introduces a different set of compactness challenges related to the nonlinearity itself.
@@
    There exists a nontrivial solution $u_{0}\in X^{1}(\mathbb{R}^{N})$ of (1.2) that satisfies the Pohožaev constraint $\mathcal{M}$} and $J(u_{0})>0$.
@@
Next, we investigate the existence of non-radial solutions. We follow the ideas of [18]. Suppose that $N\geq 4$ and write $\mathbb{R}^{N}=\mathbb{R}^{M}\times\mathbb{R}^{M}\times\mathbb{R}^{N-2M}$, where $2\leq M\leq\frac{N}{2}$. For $x\in\mathbb{R}^{N}$ we will write $x=(x_{1},x_{2},x_{3})$, where $x_{1},x_{2}\in\mathbb{R}^{M}$ and $x_{3}\in\mathbb{R}^{N-2M}$. We introduce an action $\tau:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ by
@@
Suppose that $N\geq 4$, (G1)–(G5) are satisfied and that $g$ is odd. Then
$\inf_{{\mathcal{M}}\cap X_{\tau}\cap X^{1}_{\mathcal{O}}(\mathbb{R}^{N})}J\geq 2\inf_{{\mathcal{M}}}J>0$
@@ 2. Functional setting
Let $u\in X^{1}(\mathbb{R}^{N})\setminus H^{1}(\mathbb{R}^{N})$. Note that, thanks to e.g. the characterization of the norm (2.4) from [26, 27], we know that $\nabla u\in L^{2}(\mathbb{R}^{N}\setminus B(0,\varepsilon))$ for every $\varepsilon>0$. Moreover, since $u\not\in H^{1}(\mathbb{R}^{N})$, $\nabla u\not\in L^{2}(B(0,\varepsilon))$ for every $\varepsilon>0$. Fix $y\in\mathbb{R}^{N}\setminus\{0\}$ and choose $\varepsilon<\frac{|y|}{2}$. Then $\nabla u(\cdot-y)\not\in L^{2}(B(y,\varepsilon))$, but $B(y,\varepsilon)\subset\mathbb{R}^{N}\setminus B(0,\varepsilon)$, so $\nabla u(\cdot-y)\not\in L^{2}(\mathbb{R}^{N}\setminus B(0,\varepsilon))$, meaning that $u(\cdot-y)\not\in X^{1}(\mathbb{R}^{N})$. ∎
@@ 3. Variational setting
Then, by the deformation lemma [29, Lemma 2.3], we may find a flow $\eta\in{\mathcal{C}}([0,1]\times X^{1}(\mathbb{R}^{N});X^{1}(\mathbb{R}^{N}))$ such that $\eta(1,\gamma(\cdot))\in\Gamma$ and $J(\eta(1,\gamma(t^{*})))\leq c-\varepsilon<c$. Moreover, for every $t\neq t^{*}$, $J(\eta(1,\gamma(t)))\leq J(\gamma(t))<c$. Hence, for every $t\in[0,1]$ we have $J(\eta(1,\gamma(t)))<c$. Thus $\sup_{t\in[0,1]}J(\eta(1,\gamma(t)))<c$ and we obtain a contradiction. ∎
@@
For every $\varepsilon\in\left(0,\frac{c}{2}\right)$ we may choose a path $\gamma\in\Gamma$ such that $\sup_{t\in[0,1]}(J\circ\Phi)(0,\gamma(t))\leq c+\varepsilon$. Hence, by Ekeland’s variational principle [29, Theorem 2.8], there exists $(t,u)\in\mathbb{R}\times X^{1}(\mathbb{R}^{N})$ such that
@@ 4. Profile decomposition of Pohožaev-Palais-Smale sequences
Since $w^{1}\in H^{1}(\mathbb{R}^{N})$ is a weak solution to (4.5), from [21, Theorem 2], there is $\alpha>0$ such that $|w^{1}(x)|\lesssim e^{-\alpha|x|}$. For $x\in B(0,R)$ we have $|x-y_{n}|\geq|y_{n}|-|x|\geq|y_{n}|-R\to\infty$, so one has
@@ 6.2. Regularity of the solution
and $|u|^{p-2}\in L^{N/2}(\Omega)$. Therefore, by Brezis-Kato lemma (see [23, Lemma B.3]), $u\in L^{q}_{\mathrm{loc}}(\Omega)$ for any $q<\infty$. Then, by Caldéron-Zygmund inequality (see [23, Lemma B.2]), $u\in W^{2,q}_{\mathrm{loc}}(\Omega)$ and next, by Sobolev embedding theorem, $u\in C^{1,\alpha}_{\mathrm{loc}}(\Omega)$. Since $x_{0}$ was an arbitrary point, we conclude $u\in C^{1,\alpha}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$. ∎
@@ 6.3. There is a radial solution, which is a minimizer on $\mathcal{M}$}
$\int_{\mathbb{R}^{N}}\frac{|u^{*}|^{2}}{|x|^{2}}\,dx\geq\int_{\mathbb{R}^{N}}\frac{u^{2}}{|x|^{2}}\,dx,\quad\int_{\mathbb{R}^{N}}G(u^{*})\,dx=\int_{\mathbb{R}^{N}}G(u)\,dx.$
@@ 6.4. Nonnegative solutions do not belong to $H^{1}(\mathbb{R}^{N})$
From [22, Proposition 2.1], $)u\not\in H^{1}(B(0,\widetilde{\rho}$)). ∎
@@ 7. Existence of non-radial solutions
In this section we will prove Theorem 1.3. Following [18], we recall the subspace $X_{\tau}$ given by (1.4) and the group $\mathcal{O}=\mathcal{O}(M)\times\mathcal{O}(M)\times\mathcal{O}(N-2M)$ acting on $\mathbb{R}^{N}=\mathbb{R}^{M}\times\mathbb{R}^{M}\times\mathbb{R}^{N-2M}$, $N\geq 4$. Repeating the reasoning from [18, Corollary 3.2] we see that the following concentration-compactness principle holds in the subspace $H^{s}_{\mathcal{O}}(\mathbb{R}^{N})$ of $\mathcal{O}$}-invariant functions in $H^{s}(\mathbb{R}^{N})$.
@@
Let $R>1$. Now, as in [18, Remark 4.2], we choose an odd and smooth function $\varphi:\mathbb{R}\rightarrow[-1,1]$ such that $\varphi(x)=1$ for $x\geq 1$, and $\varphi(x)=-1$ for $x\leq-1$. Then, as shown in [18, Remark 4.2], the function (with a small abuse of notation $w_{R}(x)=w_{R}(|x|)$)